OpenStudy (anonymous):

Find the antiderivative F of f that satisfies the given condition. Check your answer by comparing the graphs of f and F. f(x) = 5x^4 − 4x^5, F(0) = 6

OpenStudy (anonymous):

so you take the integral (antiderivative) but you'll get a constant. Sub zeros in for all your x's and set you y to 6 and you'll be able to solve the constant

OpenStudy (anonymous):

Your familiar with the power rule for antiderivatives?

OpenStudy (anonymous):

no, we literally just started them.

OpenStudy (anonymous):

have you covered derivatives?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so the power rule states: \[\huge{\frac{d}{dx}x^n=nx^{n-1}}\] so you drop the exponent in front and reduce the exponent by one

OpenStudy (anonymous):

antiderivative would be just the opposite. For the antiderivative you first increase the exponent by one and then 'divide' by that new exponent

OpenStudy (anonymous):

ohh okay!

OpenStudy (anonymous):

\[\huge{\int\limits_{}^{}x^n=\frac{x^{n+1}}{n+1}}\]

OpenStudy (anonymous):

You'll do that for each term, and introduce the c as a constant of integration

OpenStudy (anonymous):

thanks!